# Organizing Unstructured Image Collections using Natural Language

### v4

@@ 6 Applications
Analyzing Social Media Image Popularity: What visual traits drive popularity? To explore this, we apply $\mathcal{X}$-Cluster to 4.1k Flickr photos from the SPID dataset (Ortis et al., 2019), where popularity is measured by image view count. $\mathcal{X}$-Cluster discovered ten grouping criteria and organizes photos into semantic clusters under each. Using the grouping results, Figure 9 compares the sample popularity distributions of the Top Trending and the Top Mainstream clusters for three criteria.
@@
Findings: As shown in Figure 9, combining $\mathcal{X}$-Cluster’s grouping with popularity scores provides a direct interpretation of the visual elements that drive trends versus those that define widely uploaded images. Interestingly, we find that trending elements often contrast with mainstream ones, such as “Musical activities” vs. “Rest and relaxation” or “High-intensity expressions” vs. “Neutral emotion”. These results suggest that attention-grabbing visuals stand out due to novelty or intensity, especially in today’s short attention span era (McSpadden, 2015; Farid, 2024), underscoring $\mathcal{X}$-Cluster as a powerful tool for deep dataset analysis and understanding social behavior. For full findings and additional analysis, see App. P.2.
Confirming and Mitigating Dataset Bias: Lastly, in App. P.3, we used $\mathcal{X}$-Cluster to confirm the gender bias in 162k CelebA (Liu et al., 2015) training images and successfully trained a debiased model using the predicted groupings with DRO (Sagawa et al., 2020). Tab. 35 shows our debiased model achieves robustness comparable to B2T (Kim et al., 2024), which uses annotations for bias detection.
@@ 7 Conclusion
We introduced the OpenSMC task and proposed $\mathcal{X}$-Cluster for discovering grouping criteria and semantic substructures in an unstructured image collections, expressed in natural language. We evaluated $\mathcal{X}$-Cluster on six benchmarks and three applications, showing its ability to automatically discover valuable insights stored in visual data.
@@ Acknowledgments
This work was supported by the MUR PNRR project FAIR - Future AI Research (PE00000013), funded by NextGenerationEU. E.R. acknowledges partial support from the PRECRISIS project, funded by the EU Internal Security Fund (ISFP-2022-TFI-AG-PROTECT-02-101100539), the EU project SPRING (No. 871245), and the PRIN project LEGO-AI (Prot. 2020TA3K9N). This work was conducted within the Vision and Learning joint laboratory of FBK and UNITN. We thank OpenAI for supporting our experiments through the Researcher Access Program (0000004498). We extend our gratitude to Feng Xue for his valuable suggestions on plot creation. M.L. would like to thank Margherita Potrich for her unwavering support.
Limitations
As discussed in § 5.1 and § 5.2, $\mathcal{X}$-Cluster still has significant room for improvement. It is thus not guaranteed to uncover all valid grouping criteria or perfectly align semantic substructures with user preferences. Moreover, although designed as an automatic, assistive tool for large-scale visual data analysis, $\mathcal{X}$-Cluster requires human oversight; users bear the ultimate responsibility for correctly interpreting the generated results.
Since $\mathcal{X}$-Cluster relies on multimodal large language models (MLLMs) and large language models (LLMs), it inherits their limitations, including potential errors, biases, or hallucinations. Foundation models such as LLaVA-Next (Liu et al., 2024b), CLIP (Radford et al., 2021), and Llama series (Meta, 2024b, a) reflect biases present in their internet-scale training data (Goh et al., 2021; Schramowski et al., 2022; Kirk et al., 2021; Yamada et al., 2023; Wan et al., 2023). These biases can be challenging to identify or quantify due to the vast—and sometimes proprietary—nature of the training sets (Radford et al., 2021; OpenAI, 2022a, 2023).
Furthermore, both MLLMs and LLMs may suffer from hallucinations. For LLMs (Ji et al., 2023; Yao et al., 2023a), this means generating information that appears accurate but is not. For MLLMs (or VQA models) (Ge et al., 2024), it involves creating image descriptions that seem reasonable but do not accurately reflect the visual content.
To alleviate these limitations, we provide detailed analyses of failure cases (App. I) for our proposed $\mathcal{X}$-Cluster framework, strategies for handling hallucinated or invalid criteria (App. K), and methods to mitigate biases (App. L) in the technical appendix.
Lastly, due to the zero-shot nature of $\mathcal{X}$-Cluster, its limitations may be mitigated by incorporating improved foundation models (OpenAI, 2025), employing bias alleviation techniques (Kong et al., 2024), or applying hallucination reduction strategies (Sahoo et al., 2024; Ge et al., 2024).
Ethics Statement
The motivation behind our studies on biases in existing datasets and text-to-image (T2I) generative model outputs is to reveal and address these biases that objectively exist in the datasets and models. We emphasize that our aim is to study and mitigate these issues, and in doing so, we do not create any new biases or disturbing content. Specifically, in § 6, we use well-established benchmarks, such as CelebA (Liu et al., 2015), for our study of dataset bias, and for bias discovery in T2I generative models, we select occupation-related subjects known to be associated with biases from prior studies (Bianchi et al., 2023; Bolukbasi et al., 2016). However, we acknowledge that our methodology and findings could potentially be misused by malicious actors to promote harmful narratives or discrimination against certain groups. We strongly oppose any such misuse or misrepresentation of our work. Our research is conducted with the aim of advancing technology while prioritizing public welfare and well-being.
For the creation of our two new benchmarks, COCO-4c and Food-4c, we sourced images exclusively from the COCO-val-2017 (Lin et al., 2014) and Food-101 (Bossard et al., 2014) datasets, strictly adhering to their licensing agreements. Additionally, we utilized voluntary human annotators for proposing valid grouping criteria and creating annotations along these criteria, rather than employing annotators from crowdsourcing platforms. This decision was made to ensure sustainability, fair compensation, and high-quality work, as well as to safeguard the psychological well-being of participants. Similarly, for our user study on T2I model bias evaluation, we recruited voluntary participants via questionnaires to collect human evaluation results. The user study was conducted entirely anonymously, with participants providing informed consent. Our project, including data annotation and the user study involving human subjects, was approved by the Ethical Review Board of our university.
Appendix
@@
In App. A, we first provide a reproducibility statement to ensure the transparency and replicability of our work.
App. B expands on Related Work, covering relevant tasks and methods. App. C describes the benchmarks used in our study, including the construction of the newly proposed COCO-4c and Food-4c datasets, as well as the process for creating hard ground-truth criteria for proposer evaluation. App. D details the evaluation metrics used in this work.
App. E presents the prompts and implementation details for $\mathcal{X}$-Cluster, covering both Criteria Proposers and Semantic Groupers. App. F provides implementation specifics of the compared methods. Additional quantitative results, including evaluations of the Criteria Proposer, Semantic Grouper, and comparisons with other clustering methods, are in App. G.
For qualitative analysis, App. H presents further visualizations of predicted clusters, while App. I examines failure cases. App. J investigates the impact of multi-granularity clustering. App. K explores the effect of invalid (hallucinated) criteria on system performance, and App. L studies the influence of foundation model hallucinations and biases.
App. M analyzes the computational cost and runtime of $\mathcal{X}$-Cluster. App. N and App. O extend our analysis with studies on system sensitivity and fine-grained image collections. App. P provides additional findings, implementation details, and user study results for the three explored applications. App. Q discusses how LLMs enhance image clustering. Finally, App. R outlines potential future research directions for our proposed OpenSMC task.
@@ Appendix A Reproducibility Statement
Appendix B Additional Related Work
@@
Topic Discovery. The setting of open-ended semantic multiple clustering (OpenSMC) is also related to the field of Topic Discovery (Blei et al., 2003; Wang et al., 2009; Eklund and Forsman, 2022) in natural language processing, which aims to identify textual themes from large text corpora (e.g., documents). Our work shares motivational similarities with topic discovery because both tasks seek to find common, thematic concepts from large volumes of data. In contrast, our work focuses on discovering thematic criteria from large visual content. However, indeed, the core challenges of OpenSMC and topic discovery are highly similar: they both require systems that can concurrently reason over large volumes of data. Nevertheless, OpenSMC is an even more challenging task than topic discovery for two reasons: i) semantics are not explicitly expressed in images, whereas they are in text; ii) there is currently no vision model that can encode large sets of images and reliably reason over them. Thus, in this work, we translate images to text and use text as a proxy to elicit the large-scale reasoning capability of large language models (Meta, 2024a).
@@
Multimodal Large Language Model. Recent advancements in multimodal large language models (MLLMs) have been driven by the availability of large-scale vision-language aligned training data. The typical paradigm (Liu et al., 2024c) involves using a pre-trained large language model (LLM) (Meta, 2024b; Chiang et al., 2023; Jiang et al., 2023; Meta, 2024a) alongside a pre-trained vision encoder (Radford et al., 2021). A projector is learned to align visual inputs with the LLM in the embedding space, which enhances visual understanding by utilizing the reasoning capabilities of LLMs. Several models have achieved significant success in zero-shot image captioning and visual question answering (VQA), including BLIP-2 (Li et al., 2023a), BLIP-3 (Xue et al., 2024), Kosmos-2 (Peng et al., 2024), and the LLaVA series (Liu et al., 2024c, b; Li et al., 2025). In our proposed $\mathcal{X}$-Cluster framework, we employ MLLM primarily as a zero-shot image parser, converting visual information into text and using this text as a proxy to elicit LLMs for reasoning over large image collections and discovering grouping criteria. Additionally, we leverage the multi-image reasoning capability of LLaVA-NeXT-Interleave (Li et al., 2025) to establish a baseline image-based proposer for the OpenSMC task, while utilizing BLIP-2 with customized prompts in a VQA style (Shao et al., 2023; Zhu et al., 2024) as the image-based grouper to form semantic clusters linked to specific visual content within the images.
@@
Text-Driven Image Retrieval. Given a query text (e.g., “sofa” or “person wearing a blue T-shirt”), text-driven image retrieval methods (Karthik et al., 2024; Liu et al., 2023; Wu et al., 2023) aim to find images from an image collection that are relevant to the query. In other words, in the scenario we are considering, given the image collection and a list of text queries, one can organize images according to the text using text-driven image retrieval techniques. In this context, the query can be considered as a sort of “cluster name”. However, this differs significantly from the proposed task of open-ended semantic multiple clustering (OpenSMC), because OpenSMC requires both discovering the textual criteria and the corresponding textual clusters. Thus, without knowing text queries as prior information, text-driven image retrieval methods are not able to accomplish OpenSMC.
Appendix C Benchmark Details
C.1 Construction of COCO-4c and Food-4c
@@
(1) Criteria Identification: We first split COCO-val-2017 (Lin et al., 2014) and Food-101 (Bossard et al., 2014) images into batches of 100. Each batch was stitched into a 10$\times$10 grid to form a single image. These grid images were then distributed to 5 human annotators, who were tasked with identifying grouping criteria. For each dataset, we selected the 4 most frequently occurring criteria, as shown in Tab. 3, to proceed with per-image annotation.
(2) Label Candidate Generation: To facilitate the annotation process, we used GPT-4V (OpenAI, 2023) to generate an initial list of candidate labels for each criterion. Specifically, for each criterion of COCO-4c and Food-4c, GPT-4V was prompted to assign a label that reflected the criterion for each image. This resulted in a list of criterion-specific label candidates for each dataset.
@@
[…]
Following these steps, we constructed COCO-4c and Food-4c. Note that we used the official COCO-val-2017 (Lin et al., 2014) and Food-101 (Bossard et al., 2014) images for our benchmarks and did not collect any new images. We adhered strictly to the licenses of the datasets during their creation. The exact number of classes is presented in Tab. 3. Additionally, the annotated class names for each criterion of COCO-4c are provided in Tab. 4, and for Food-4c in Tab. 5.
@@
C.2 Details on Hard Grouping Criteria Annotation
@@
In Tab. 6, we present the additional annotated Hard grouping criteria ground truth alongside the Basic criteria for each benchmark.
@@
(1) Criteria Discovery: We divided each dataset into batches of 100 images, displaying each batch in a 10$\times$10 grid. Five human annotators were assigned to each batch and instructed to identify as many valid grouping criteria as possible. The proposed criteria from each annotator were then combined to form a comprehensive set of grouping criteria for the dataset.
@@
By following this process, we developed a more comprehensive grouping criteria set as the Hard ground-truth for each benchmark, as shown in Tab. 6. This resulted in sets containing 8 criteria for Fruit-2c, 4 criteria for card, 11 criteria for Action-3c, 7 criteria for Clevr-4c, 17 criteria for COCO-4c, and 11 criteria for Food-4c. These expanded ground-truth sets enable us to more effectively evaluate the capabilities of various criteria discovery methods, providing a clearer understanding of different criteria proposers.
Appendix D Further Details of Evaluation Protocol
Further Discussion on Clustering Accuracy (CAcc). Clustering Accuracy (CAcc) (Han et al., 2021) is evaluated by applying the Hungarian algorithm (Kuhn, 1955) to determine the optimal assignment between the predicted cluster indices and ground-truth labels. As extensively discussed in the GCD (Vaze et al., 2022) literature, if the number of predicted clusters (groups) exceeds the total number of ground-truth classes (groups), the extra clusters (not matched by the Hungarian algorithm) are assigned to a null set, and all instances in those clusters are considered incorrect during evaluation. On the other hand, if the number of predicted clusters is lower than the number of ground-truth classes, the extra classes are assigned to a null set, and all instances with those ground-truth labels are similarly considered incorrect. Thus, CAcc is maximized only when the number of predicted clusters matches the number of ground-truth clusters.

### v5

@@ 6.2 Analyzing Social Media Image Popularity
What makes a photo popular? To explore this, we apply $\mathcal{X}$-Cluster to 4.1k Flickr photos from the SPID dataset [ortis2019prediction], where popularity is measured by image view count. $\mathcal{X}$-Cluster discovered 10 grouping criteria and organizes photos into semantic clusters under each. Using the grouping results, Figure 9 compares the sample popularity distributions of the Top Trending and the Top Mainstream clusters across three criteria.
@@
Findings: As shown in Figure 9, combining $\mathcal{X}$-Cluster’s grouping with popularity scores provides a direct interpretation of the visual elements that drive trends versus those that define widely uploaded images. Interestingly, we find that trending elements often contrast with mainstream ones, such as “Musical activities” vs. “Rest and relaxation” or “High-intensity expressions” vs. “Neutral emotion”. These results suggest that attention-grabbing visuals stand out due to novelty or intensity, especially in today’s short attention span era [mcspadden2015you, farid_2024], underscoring $\mathcal{X}$-Cluster as a powerful tool for deep dataset analysis and understanding social behavior. For full findings and additional analysis, see Supp. Q.2.
@@ 7 Conclusion
We introduce the OpenSMC task and propose $\mathcal{X}$-Cluster, a system that discovers interpretable grouping criteria and substructures in image collections, effectively extracting valuable insights across six datasets and three applications.
@@ Acknowledgments
This work was supported by the EU Horizon projects ELIAS (No. 101120237) and ELLIOT (No. 101214398). We greatly acknowledge CINECA and the ISCRA initiative for providing high-performance computing resources. We also extend our gratitude to Feng Xue for his valuable suggestions on plot creation. M.L. would like to thank Margherita Potrich for her unwavering support.
@@
In Supp. A, we provide a reproducibility statement to ensure the transparency and replicability of our work. Additionally, in Supp. B, we present an ethics statement to discuss and address the potential ethical implications and concerns associated with the proposed task and methodology, highlighting possible societal impacts and mitigation strategies.
Supp. C expands on Related Work, covering relevant tasks and methods. Supp. D describes the benchmarks used in our study, including the construction of the newly proposed COCO-4c and Food-4c datasets, as well as the process for creating hard ground-truth criteria for proposer evaluation. Supp. E details the evaluation metrics used in this work.
Supp. F presents the prompts and implementation details for $\mathcal{X}$-Cluster, covering both Criteria Proposers and Semantic Groupers. Supp. G provides implementation specifics of the compared methods. Additional quantitative results, including evaluations of the Criteria Proposer, Semantic Grouper, and comparisons with other clustering methods, are in Supp. H.
For qualitative analysis, Supp. I presents further visualizations of predicted clusters, while Supp. J examines failure cases. Supp. K investigates the impact of multi-granularity clustering. Supp. L explores the effect of invalid (hallucinated) criteria on system performance, and Supp. M studies the influence of foundation model hallucinations and biases.
Supp. N analyzes the computational cost and runtime of $\mathcal{X}$-Cluster. Supp. O and Supp. P extend our analysis with studies on system sensitivity and fine-grained image collections. Supp. Q provides additional findings, implementation details, and user study results for the three explored applications. Supp. R discusses how LLMs enhance image clustering. Finally, Supp. S outlines potential future research directions for our proposed OpenSMC task.
@@ Appendix A Reproducibility Statement
Appendix B Ethics Statement
The motivation behind our studies on biases in existing datasets and text-to-image (T2I) generative model outputs is to reveal and address these biases that objectively exist in the datasets and models. We emphasize that our aim is to study and mitigate these issues, and in doing so, we do not create any new biases or disturbing content. Specifically, in Sec. 6, we use well-established benchmarks, such as CelebA [liu2015deep], for our study of dataset bias, and for bias discovery in T2I generative models, we select occupation-related subjects known to be associated with biases from prior studies [bianchi2023easily, bolukbasi2016man]. However, we acknowledge that our methodology and findings could potentially be misused by malicious actors to promote harmful narratives or discrimination against certain groups. We strongly oppose any such misuse or misrepresentation of our work. Our research is conducted with the aim of advancing technology while prioritizing public welfare and well-being.
For the creation of our two new benchmarks, COCO-4c and Food-4c, we sourced images exclusively from the COCO-val-2017 [lin2014microsoft] and Food-101 [bossard2014food] datasets, strictly adhering to their licensing agreements. Additionally, we utilized voluntary human annotators for proposing valid grouping criteria and creating annotations along these criteria, rather than employing annotators from crowdsourcing platforms. This decision was made to ensure sustainability, fair compensation, and high-quality work, as well as to safeguard the psychological well-being of participants. Similarly, for our user study on T2I model bias evaluation, we recruited voluntary participants via questionnaires to collect human evaluation results. The user study was conducted entirely anonymously, with participants providing informed consent. Our project, including data annotation and the user study involving human subjects, was approved by the Ethical Review Board of our university.
Appendix C Additional Related Work
@@
Topic Discovery. The setting of open-ended semantic multiple clustering (OpenSMC) is also related to the field of Topic Discovery [blei2003latent, wang2009multi, eklund2022topic] in natural language processing, which aims to identify textual themes from large text corpora (e.g., documents). Our work shares motivational similarities with topic discovery because both tasks seek to find common, thematic concepts from large volumes of data. In contrast, our work focuses on discovering thematic criteria from large visual content. However, indeed, the core challenges of OpenSMC and topic discovery are highly similar: they both require systems that can concurrently reason over large volumes of data. Nevertheless, OpenSMC is an even more challenging task than topic discovery for two reasons: i) semantics are not explicitly expressed in images, whereas they are in text; ii) there is currently no vision model that can encode large sets of images and reliably reason over them. Thus, in this work, we translate images to text and use text as a proxy to elicit the large-scale reasoning capability of large language models [meta2024llama31].
@@
Multimodal Large Language Model. Recent advancements in multimodal large language models (MLLMs) have been driven by the availability of large-scale vision-language aligned training data. The typical paradigm [liu2024visual] involves using a pre-trained large language model (LLM) [meta2024llama3, vicuna2023, jiang2023mistral, meta2024llama31] alongside a pre-trained vision encoder [radford2021learning]. A projector is learned to align visual inputs with the LLM in the embedding space, which enhances visual understanding by utilizing the reasoning capabilities of LLMs. Several models have achieved significant success in zero-shot image captioning and visual question answering (VQA), including BLIP-2 [li2023blip], BLIP-3 [xue2024xgen], Kosmos-2 [peng2023kosmos], and the LLaVA series [liu2024visual, liu2024llava, li2024llava]. In our proposed $\mathcal{X}$-Cluster framework, we employ MLLM primarily as a zero-shot image parser, converting visual information into text and using this text as a proxy to elicit LLMs for reasoning over large image collections and discovering grouping criteria. Additionally, we leverage the multi-image reasoning capability of LLaVA-NeXT-Interleave [li2024llava] to establish a baseline image-based proposer for the OpenSMC task, while utilizing BLIP-2 with customized prompts in a VQA style [shao2023prompting, zhu2023chatgpt] as the image-based grouper to form semantic clusters linked to specific visual content within the images.
@@
Text-Driven Image Retrieval. Given a query text (e.g., “sofa” or “person wearing a blue T-shirt”), text-driven image retrieval methods [karthik2024visionbylanguage, liu2023zero, wu2023cap4video] aim to find images from an image collection that are relevant to the query. In other words, in the scenario we are considering, given the image collection and a list of text queries, one can organize images according to the text using text-driven image retrieval techniques. In this context, the query can be considered as a sort of “cluster name”. However, this differs significantly from the proposed task of open-ended semantic multiple clustering (OpenSMC), because OpenSMC requires both discovering the textual criteria and the corresponding textual clusters. Thus, without knowing text queries as prior information, text-driven image retrieval methods are not able to accomplish OpenSMC.
Appendix D Benchmark Details
D.1 Construction of COCO-4c and Food-4c
@@
(1) Criteria Identification: We first split COCO-val-2017 [lin2014microsoft] and Food-101 [bossard2014food] images into batches of 100. Each batch was stitched into a $\times10$ grid to form a single image. These grid images were then distributed to 5 human annotators, who were tasked with identifying grouping criteria. For each dataset, we selected the 4 most frequently occurring criteria, as shown in Tab. 1, to proceed with per-image annotation.
(2) Label Candidate Generation: To facilitate the annotation process, we used GPT-4V [openai2023gpt4] to generate an initial list of candidate labels for each criterion. Specifically, for each criterion of COCO-4c and Food-4c, GPT-4V was prompted to assign a label that reflected the criterion for each image. This resulted in a list of criterion-specific label candidates for each dataset.
@@
(4) Label Merging: Image annotation is inherently subjective, with annotators potentially assigning different labels for the same criterion. For example, one annotator might label the Mood criterion as “Happy”, while another might label it as “Joyful” or “Delightful”. To resolve such discrepancies, we used majority voting to determine the final label for each image. Specifically, the most frequently assigned label among the 10 annotators was chosen as the final label for each criterion.
Following these steps, we constructed COCO-4c and Food-4c. Note that we used the official COCO-val-2017 [lin2014microsoft] and Food-101 [bossard2014food] images for our benchmarks and did not collect any new images. We adhered strictly to the licenses of the datasets during their creation. The exact number of classes is presented in Tab. 1. Additionally, the annotated class names for each criterion of COCO-4c are provided in Tab. 2, and for Food-4c in Tab. 3.
@@
D.2 Details on Hard Grouping Criteria Annotation
@@
In § D.2, we present the additional annotated Hard grouping criteria ground truth alongside the Basic criteria for each benchmark.
@@
(1) Criteria Discovery: We divided each dataset into batches of 100 images, displaying each batch in a $\times10$ grid. Five human annotators were assigned to each batch and instructed to identify as many valid grouping criteria as possible. The proposed criteria from each annotator were then combined to form a comprehensive set of grouping criteria for the dataset.
@@
By following this process, we developed a more comprehensive grouping criteria set as the Hard ground-truth for each benchmark, as shown in § D.2. This resulted in sets containing 8 criteria for Fruit-2c, 4 criteria for card, 11 criteria for Action-3c, 7 criteria for Clevr-4c, 17 criteria for COCO-4c, and 11 criteria for Food-4c. These expanded ground-truth sets enable us to more effectively evaluate the capabilities of various criteria discovery methods, providing a clearer understanding of different criteria proposers.
Appendix E Further Details of Evaluation Protocol
Further Discussion on Clustering Accuracy (CAcc). Clustering Accuracy (CAcc) [han2021autonovel] is evaluated by applying the Hungarian algorithm [kuhn1955hungarian] to determine the optimal assignment between the predicted cluster indices and ground-truth labels. As extensively discussed in the GCD [vaze2022generalized] literature, if the number of predicted clusters (groups) exceeds the total number of ground-truth classes (groups), the extra clusters (not matched by the Hungarian algorithm) are assigned to a null set, and all instances in those clusters are considered incorrect during evaluation. On the other hand, if the number of predicted clusters is lower than the number of ground-truth classes, the extra classes are assigned to a null set, and all instances with those ground-truth labels are similarly considered incorrect. Thus, CAcc is maximized only when the number of predicted clusters matches the number of ground-truth clusters.